OpenStudy (anonymous):

suppose f(x) = ax+b where a and be are constants. if f(1)=7 and f(3)=5, find a and b.

OpenStudy (solomonzelman):

Ok, \(f(1)=7\) means that \(\color{#000000 }{ \displaystyle f(1) = a\cdot (1)+b=7 }\) \(\color{#000000 }{ \displaystyle a+b=7 }\)

OpenStudy (anonymous):

ok...i kinda get it

OpenStudy (solomonzelman):

plug in f(3)=5 into the function (as I did with f(1)=7) and you will have the second equation

OpenStudy (anonymous):

ok so then it becomes \[f(3)=a(3)+b=-5\]

OpenStudy (anonymous):

am i correct or..?

OpenStudy (solomonzelman):

-5 ?

OpenStudy (anonymous):

sorry, the f(3)=-5

OpenStudy (anonymous):

i made a typo up above

OpenStudy (solomonzelman):

I thought the 5 was positive

OpenStudy (solomonzelman):

oh, typo, ok

OpenStudy (anonymous):

yea haha sorry. so do we leave the equations like so or do we do more with them

OpenStudy (solomonzelman):

yes, and a(3), is same thing as a•3, so you can just write \(\color{#000000 }{ \displaystyle 3a+b=-5 }\)

OpenStudy (solomonzelman):

So you were given f(1)=7 and f(3)=-5 and thanks to that we have the two equations below, \(\color{#000000 }{ \displaystyle a+b=7 }\) \(\color{#000000 }{ \displaystyle 3a+b=-5 }\) which is just a system of equations.

OpenStudy (anonymous):

the back of my book says the answers are a=-6 and b=13

OpenStudy (solomonzelman):

and you want to know why...

OpenStudy (solomonzelman):

*(duh)*

OpenStudy (anonymous):

yes please lol

OpenStudy (solomonzelman):

first, is it clear how we obtained the system of equations below? \(\color{#000000 }{ \displaystyle a+b=7 }\) \(\color{#000000 }{ \displaystyle 3a+b=-5 }\)

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

You can subtract b on both sides of the first equation, and what wll this give you?

OpenStudy (anonymous):

um you get \[a=7-b\]

OpenStudy (anonymous):

i think

OpenStudy (solomonzelman):

yes, and substitute this into the second equation for a \(\color{#000000 }{ \displaystyle 3(7-b)+b=-5 }\)

OpenStudy (solomonzelman):

then expand and simplify the left side

OpenStudy (anonymous):

ok so then you get \[21-2b=-5\]

OpenStudy (anonymous):

then subtract 21 from 5, you get -16/2b

OpenStudy (anonymous):

which is -8

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

for b?

OpenStudy (anonymous):

wait no i got it wrong

OpenStudy (anonymous):

oh wait haha i see what i did wrong, ok b=13

OpenStudy (anonymous):

now that i know that, how do i get a?

OpenStudy (anonymous):

oh wait you would just subtract 7-13=-6

OpenStudy (anonymous):

ok hahaha makes so much sense now

OpenStudy (solomonzelman):

21−2b=−5 −2b=−5-21

OpenStudy (solomonzelman):

-26

OpenStudy (solomonzelman):

not -16

OpenStudy (anonymous):

yea haha i see what i did wrong

OpenStudy (solomonzelman):

and b is?

OpenStudy (anonymous):

13

OpenStudy (solomonzelman):

as given in answers...

OpenStudy (anonymous):

b is 13 in the back of the book...

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

and then the a, a+b=7 a+13=7 a=...

OpenStudy (anonymous):

great! ok if u dont mind me asking, how long will you be online for? a is -6

OpenStudy (anonymous):

i just wanna know just in case i accidentally bother you lol

OpenStudy (solomonzelman):

No, it is fine... we all make simple error. But when teachers take off 10 points off for that, they are pigs !

OpenStudy (solomonzelman):

yw, and good luck

OpenStudy (anonymous):

haha thanks